OpenStudy (anonymous):

Solve the equation: (3x - 2)(7x + 1) = 0.

Parth (parthkohli):

There's a rule that you use here - the Zero Product Rule(tm).

Parth (parthkohli):

If you have two factors \(a,b\) - and \(a\times b = 0\), then either a = 0 or b = 0.

OpenStudy (jiteshmeghwal9):



Parth (parthkohli):

Using the same rule, we have \[3x - 2 = 0\\ OR\\7x + 1 = 0 \]When you solve these two equations, you'd get the roots.

OpenStudy (anonymous):

so i just add and subtract those so 3x-2 becomes x and 7x+1 becomes 8x?

OpenStudy (anonymous):